Tickets for a concert were sold to adults for $3 and students for $2 if the total receipts were $824 and twice as many adult tickets as students were sold then how many of each were sold?
Answer by nerdybill(7384)

Let x = number of student tickets sold
the because "twice as many adult tickets as students were sold"
2x = number of adult tickets sold
.
2x + 3(2x) = 824
2x + 6x = 824
8x = 824
x = 103 (number of student tickets
.
Adult tickets:
2x = 2(103) = 206 (number of adult tickets)